# How to Make the Most of LLMs’ Grammatical Knowledge for Acceptability Judgments

### v1

@@ C.1 Max Accuracy
Llama-3 Llama-3 -Instruct Mixtral Mixtral -Instruct Qwen2 Qwen2 -Instruct Qwen2 Qwen2 -Instruct
LP 79.6 77.1 82.5 82.3 80.4 79.7 85.4 85.4
MeanLP 77.1 74.8 79.6 79.4 77.7 77.1 74.5 74.3
PenLP 79.2 76.8 82.2 82.0 79.9 79.2 82.2 82.0
In-template LP $\mathbf{85.0}$ 84.2 84.6 84.5 84.1 81.3 $\mathbf{88.2}$ $\mathbf{86.5}$
In-template MeanLP 83.1 82.6 83.2 83.1 82.8 80.5 78.9 79.0
In-template PenLP 84.4 83.5 84.5 84.3 83.6 81.2 83.8 83.5
In-template comparative LP 76.4 66.0 75.2 69.8 67.7 63.6 74.2 63.5
A/B prompting 79.5 83.9 81.9 83.6 81.5 82.8 83.2 82.5
Yes/No prob comp 77.8 $\mathbf{89.3}$ $\mathbf{85.6}$ $\mathbf{87.5}$ $\mathbf{89.2}$ $\mathbf{87.4}$ 86.6 85.2
Table 9: Percentage max accuracy by method and model.
C.2 Why A/B prompting does not perform well
BLiMP CLiMP
Llama-3 Llama-3-Instruct Mixtral Mixtral-Instruct Qwen2 Qwen2-Instruct Qwen2 Qwen2-Instruct
55.0 56.6 70.1 45.9 54.0 45.7 35.9 46.1
@@ C.2 Why A/B prompting does not perform well
The low performance of A/B prompting can be partly attributed to a preference for a specific choice, A or B. Table 10 shows all our models are at least 7 points more likely to predict one of the choices over the other one, even though the gold label is sampled from a uniform distribution. This suggests that the current LLMs suffer from selection bias on multiple choices as argued by Zheng et al. (2024).
C.3 Is Yes/No prob comp strong where the acceptable sentence is longer than the unacceptable?
@@
Figure 6 shows that Yes/No prob comp is not stronger than in-template LP in phenomena where the acceptable sentence is, on average, longer than the unacceptable one. We only find no or weak negative correlations between the accuracy difference and token-length difference.

### v2

@@ D.2 Why A/B prompting does not perform well
Llama-3 Llama-3 -Inst. Mixtral Mixtral -Inst. Qwen2 Qwen2 -Inst. Llama-3 Llama-3 -Inst. Yi-1.5 Yi-1.5 -Chat Qwen2 Qwen2 -Inst.
55.0 56.6 70.1 45.9 54.0 45.7 70.5 72.8 52.7 46.7 35.9 46.1
@@
The low performance of A/B prompting can be partly attributed to a preference for a specific choice identifier, A or B. Table 12 shows all models except for Yi-1.5 models are at least 7 points more likely to predict one choice over the other one, even though the gold label is sampled from a uniform distribution. This suggests that the current LLMs suffer from selection bias on multiple choices as argued by Zheng et al. (2024).
D.3 Is Yes/No prob comp strong where the acceptable sentence is longer than the unacceptable?
@@
Figure 6 shows that Yes/No prob comp is not particularly accurate compared to in-template LP in phenomena where the acceptable sentence is, on average, longer than the unacceptable one. We only find no or weak negative correlations between the accuracy difference and token-length difference.
D.4 Voting Ensembles
Llama-3 Llama-3-Instruct Mixtral Mixtral-Instruct Qwen2 Qwen2-Instruct
Ensemble P-only 76.0 89.0 85.4 84.4 89.5 87.0
Ensemble Mix-P3 82.7 $\mathbf{89.7}$ $\mathbf{87.5}$ $\mathbf{86.6}$ $\mathbf{90.2}$ $\mathbf{87.7}$
Ensemble Mix-L3 $\mathbf{86.1}$ 86.3 86.5 86.4 86.7 84.8
Ensemble L-only 85.1 84.1 84.6 84.3 84.3 81.0
In-template LP (oracle) 85.0 84.2 84.6 84.5 84.1 81.3
Yes/No prob comp (oracle) 77.8 89.3 85.6 87.5 89.2 87.4
(a) BLiMP
Llama-3 Llama-3-Instruct Yi-1.5 Yi-1.5-Chat Qwen2 Qwen2-Instruct
Ensemble P-only 76.0 76.6 78.4 81.9 87.6 84.1
Ensemble Mix-P3 79.1 78.4 79.9 83.9 89.4 87.1
Ensemble Mix-L3 $\mathbf{87.1}$ $\mathbf{83.8}$ $\mathbf{88.7}$ $\mathbf{88.9}$ $\mathbf{90.5}$ $\mathbf{89.4}$
Ensemble L-only 86.1 83.0 88.0 87.7 88.4 87.0
In-template LP (oracle) 86.1 83.2 87.9 87.8 88.2 86.5
Yes/No prob comp (oracle) 78.1 78.4 79.7 82.1 87.5 84.3
(b) CLiMP